# Deep Unfolding Network with Spatial Alignment for multi-modal MRI reconstruction

### v1

@@ 5.3 Results on the in-house dataset
Fig. 7: Comparison on the learning trajectories of different models on the fastMRI dataset for 4x acceleration under equispaced mask.
@@
Quantitative results on the in-house dataset for 1D equispaced and random masks under 4× and 8× acceleration are presented in Table 5 and Table 6, showing that the proposed DUN-SA outperforms other methods in achieving the best results for every metric across all settings.
Qualitative results on the in-house dataset for 1D equispaced mask under 8× acceleration are depicted in Fig. 6. Similarly, by comparing the reconstructed images of different methods and their corresponding error maps, it can be observed that the proposed DUN-SA obtains the best results, reconstructing more details. Specifically, in zoomed-in regions of interest in the second row, the details indicated by the red arrow and green row are only reconstructed by the proposed DUN-SA. This in line with the performance improvement observed in the aforementioned two datasets and indicates the promising generalizability of DUN-SA.
@@
6 Discussion
In this section, we first compare the convergence performance of the proposed DUN-SA with other methods during the training process. Then, we conduct experiments to evaluate the performance under different scales of misalignment. We further compare its spatial alignment performance under different accelerations and conduct ablation studies to validate the benefits of the key components of the DUN-SA at last.
6.1 Convergence of DUN-SA vs. other methods
In this section, we compare the convergence performance of the proposed DUN-SA with other methods by recording and comparing the reconstruction performance of each method throughout the training process. Specifically, we record the PSNR, SSIM, and MAE metrics during the training process on the fastMRI dataset for 4x acceleration under equispaced mask. The curves in Fig. 7 illustrate that the convergence performance of DUN-SA is superior compared to the other methods because it consistently maintains higher PSNR and SSIM values and a lower MAE throughout the entire training process.
@@ 6.2 Performance evaluation under different scales of misalignment between modalities
To demonstrate the superior robustness of the proposed DUN-SA to misalignment between modalities, we compare it with SAN, another reconstruction method that considers misalignment. We perform different scales of spatial misalignment on the reference modality and then test using the parameters in Section 5.1. This experiment is conducted on the fastMRI dataset for 4x and 8x acceleration under both equispaced and random 1D subsampling masks. The method for simulating spatial misalignment is consistent with the approach in SAN; specifically, we employ random rotations within the range of $[-0.01\pi\sigma,0.01\pi\sigma]$, translations between $[-0.05N\sigma,0.05N\sigma]$, and displacement field bicubically interpolated from $9\times 9$ control-points (with displacements uniformly sampled within $[-0.02N\sigma,0.02N\sigma]$ in both directions. Here, $N$ represents the size of the MR images and $\sigma$ indicates the factor controlling the degree of spatial misalignment). Fig. 8 compares the reconstruction performance of the two methods under different degrees of spatial misalignment. It’s noteworthy that as the scale of spatial misalignment increases, the performance of both methods declines. However, DUN-SA degrades more slowly, whereas the performance of SAN decreases relatively faster, leading to a gradually increasing difference between the two methods. This highlights the greater robustness of the proposed DUN-SA to spatial misalignment.
6.3 Comparison of spatial alignment under different acceleration factors
Fig. 10: Quantitative comparison of MRI reconstruction (in PSNR/SSIM) with proposed DUN-SA lacking of different key components on the fastMRI dataset for 4x and 8x acceleration under equispaced and random subsampling masks. In each box plot, the average value is marked with the red diamond symbol.
@@ 6.3 Comparison of spatial alignment under different acceleration factors
In this section, we evaluate the spatial alignment performance of the proposed DUN-SA under different acceleration factors and compare it with SAN. In Fig. 9, (a) and (b) represent fully-sampled T2-weighted image and fully-sampled T1-weighted image, respectively. (c), (e), and (g) depict T1-weighted images aligned using SAN under different acceleration factors, while (d), (f), and (h) display T1-weighted images aligned using DUN-SA under different acceleration factors. We use a grid to facilitate the observation of spatial positions and employ a checkerboard to visualize misalignment. We find that both DUN-SA and SAN exhibit good spatial alignment performance at 4x acceleration. However, as the acceleration factor increases, SAN gradually fails to align T1 and T2 well, whereas DUN-SA is minimally affected. For instance, we can observe in the zoomed-in views of the third and fifth rows, at the locations indicated by the red, blue, and green arrows, misalignment still exists in (e) and (g), but this misalignment has been alleviated in (f) and (h). This demonstrates that DUN-SA, by iteratively solving the proposed model, appropriately integrates the spatial alignment task into the reconstruction process and achieves more precise spatial alignment results under high acceleration factors.
@@ 6.4 Ablation Study
To determine the optimal network architecture, we conduct two ablation studies. The first study focus on the number of stages, and the second emphasize each component within the network.
@@ 6.4.1 Effect of number of stages
To demonstrate how the number of stages $k$ affects the reconstruction performance, we carry out a quantitative comparison of DUN-SA under different numbers of stages on the fastMRI dataset for 4x and 8x acceleration under equispaced and random subsampling masks. Fig. 11 shows the mean SSIM and PSNR values for stages ranging from 1 to 16. We observed that as the number of stages increases, the quality of reconstruction improves across all settings. However, when $k=12$, further increasing the stages leads to only subtle improvements in reconstruction performance. Taking both performance and model complexity into consideration, we chose the model with $k=12$.
Fig. 11: The PSNR and SSIM curves on the fastMRI dataset with different numbers of stages k
@@ 6.4.2 Effect of key components
Table 7: Effect of each component on the performance of DUN-SA on the fastMRI dataset for 4x and 8x acceleration under equispaced and random subsampling masks, measured in PSNR and SSIM.
Methods Equispaced 4x Equispaced 8x Random 4x Random 8x
Zero-filling 26.92/0.7321 24.36/0.6428 27.12/0.7373 23.88/0.5818
w/o MM 39.36/0.9779 37.32/0.9694 43.35/0.9875 36.37/0.9645
w/o SA 41.07/0.9829 39.75/0.9787 44.68/0.9899 38.65/0.9755
w/o DB 41.16/0.9831 39.86/0.9791 44.70/0.9900 38.75/0.9758
DUN-SA 41.48/0.9838 40.23/0.9802 45.06/0.9907 39.23/0.9770
Spatial Alignment Module (SAM): The ablation study for SAM involves removing the SAM module from the network. In this scenario, the spatial alignment operator is ignored within the network. This configuration is termed the proposed DUN-SA without spatial alignment (w/o SA).
@@
Spatial Alignment Augmented Multi-Modal Learning Block (SAAMMLB): The ablation study for SAAMMLB means removing the SAAMMLB module from the network. It’s worth noting that, when the SAAMMLB module is removed, the assistance from the reference modality will also be omitted, and the spatial alignment operator will be simultaneously disregarded. This configuration is termed the proposed DUN-SA without multi-modality (w/o MM).
@@
Denoising Block (DB): The ablation study for DB refers to the removal of the DB module from the network. This configuration is termed the proposed DUN-SA without denoising block (w/o DB).
Firstly, Fig. 10 and Table 7 quantitatively demonstrate the superiority of our model. In Fig. 10, we use a box plot to depict the performance of each method. The orange horizontal line represents the median, and the red diamond represents the mean. From left to right, they are w/o MM, w/o SA, w/o DB and our proposed model (DUN-SA). It can be seen from the Fig. 10 and Table 7 that the reconstruction results on the fastMRI dataset using a single modality (w/o MM) are the poorest across all settings; the reconstruction performance decreases without the spatial alignment, indicating that after aligning the reference modality image, it can provide richer inter-modality prior to assist in the reconstruction task. Similarly, the absence of the denoising block (w/o DB) results in a noticeable decline in performance. Overall, each key component can help the model enhance reconstruction performance, and the optimal model can be obtained by combining all the modules.

### v2

@@ 5.3 Results on the In-house dataset
Quantitative results on the In-house dataset for 1D equispaced and random masks under 4× and 8× acceleration are presented in Table 3, showing that the proposed DUN-SA outperforms other methods in achieving the best results for every metric across all settings.
Qualitative results on the In-house dataset for 1D equispaced mask under 8× acceleration are depicted in Fig. 6. Similarly, by comparing the reconstructed images of different methods and their corresponding error maps, it can be observed that the proposed DUN-SA obtains the best results, reconstructing more details. Specifically, in zoomed-in regions of interest shown in the second row, the details indicated by the red arrow and green row are only reconstructed by the proposed DUN-SA. This in line with the performance improvement observed in the aforementioned two datasets and indicates the promising generalizability of DUN-SA.
5.4 Results on the BraTS 2018 dataset
Fig. 8: Comparison on the learning trajectories of different models on the fastMRI dataset for 4$\times$ acceleration under equispaced mask.
@@ 5.4 Results on the BraTS 2018 dataset
In this section, we conduct experiments on the BraTS 2018 dataset to assess model performance with greater variability in protocols and pathologies. Specifically, we test the cases where the Reference/Target modalities are T2/FLAIR; T1/FLAIR; T1CE/FLAIR; T1/T1CE; T1CE/T2. Moreover, we compare the selected best-performing four multi-modal methods across the three mentioned datasets.
Quantitative results on the BraTS 2018 dataset for 1D equispaced and random masks under 4$\times$ and 8$\times$ acceleration are presented in Table 4. The results indicate that the proposed DUN-SA maintains its high performance when facing variability in protocols and pathologies, further illustrating the generalizability of DUN-SA.
Qualitative results on the BraTS 2018 dataset for 1D random mask under 8× acceleration are depicted in Fig. 7. It can be observed from the Fig. 7 that, even when reconstructing MRI images with tumor, DUN-SA still recovers the most details, maintaining the best performance. Especially, in the zoomed-in regions of interest in the second row, we demonstrate the reconstruction performance around the tumor. In the areas indicated by green arrows, only DUN-SA and MC-CDic are able to reconstruct such signals; however, the signals reconstructed by MC-CDic are comparatively vaguer. Furthermore, within the yellow circles, DUN-SA recovered more precise and clearer edge details.
6 Discussion
In this section, we first compare the convergence performance of our proposed DUN-SA method with other methods during the training process. Subsequently, we carry out experiments to assess its performance under varying degrees of misalignment. Furthermore, we compare its spatial alignment capabilities with those of alternative methods. We evaluate the adaptability of the reconstruction performance in scenarios where the reference data is imperfect. Ablation studies are conducted to determine the optimal number of stages and to validate the advantages of the key components. We employ DUN-SA to execute a model verification experiment aimed at elucidating the operational mechanism behind the network modules. Finally, we analyze the model complexity.
6.1 Convergence of DUN-SA vs. other methods
In this section, we compare the convergence performance of the proposed DUN-SA with other methods by recording and comparing the reconstruction performance of each method throughout the training process. Specifically, we record the PSNR, SSIM, and MAE metrics during the training process on the fastMRI dataset for 4$\times$ acceleration under equispaced mask. The curves in Fig. 8 illustrate that the convergence performance of DUN-SA is superior compared to the other methods because it consistently maintains higher PSNR and SSIM values and a lower MAE throughout the entire training process.
Fig. 10: The visualization of the effects of spatial alignment on the fastMRI dataset. (a) shows the original fully-sampled T1 image. (b) represents the results of aligning an under-sampled T2 image with a fully-sampled T1 image by traditional method. (c) depicts the results of integrating traditional spatial alignment with reconstruction for joint optimization. (d) displays the result of DUN-SA. Details are shown in the first row: zoomed-in view of aligned T1 images and third row: zoomed-in views of checkerboard visualizations.
@@ 6.2 Performance evaluation under different scales of misalignment between modalities
To demonstrate the superior robustness of the proposed DUN-SA to misalignment between modalities, we compare it with SAN, another reconstruction method that considers misalignment. We perform different scales of spatial misalignment on the reference modality and then test using the parameters in Section 5.1. This experiment is conducted on the fastMRI dataset for 4$\times$ and 8$\times$ acceleration under both equispaced and random 1D subsampling masks. The method for simulating spatial misalignment is consistent with the approach in SAN; specifically, we employ random rotations within the range of $[-0.01\pi\sigma,0.01\pi\sigma]$, translations between $[-0.05N\sigma,0.05N\sigma]$, and displacement field bicubically interpolated from $9\times 9$ control-points (with displacements uniformly sampled within $[-0.02N\sigma,0.02N\sigma]$ in both directions. Here, $N$ represents the size of the MR images and $\sigma$ indicates the factor controlling the degree of spatial misalignment). Fig. 9 compares the reconstruction performance of the two methods under different degrees of spatial misalignment. It’s noteworthy that as the scale of spatial misalignment increases, the performance of both methods declines. However, DUN-SA degrades more slowly, whereas the performance of SAN decreases relatively faster, leading to a gradually increasing difference between the two methods. This highlights the greater robustness of the proposed DUN-SA to spatial misalignment.
6.3 Comparison of spatial alignment performance
In this section, we compare our spatial alignment result with an effective and widely-used traditional spatial alignment method [25]. In Fig. 10, (a) shows the original fully-sampled T1 image, demonstrating the misalignment between modalities; (b) represents the results of aligning an under-sampled T2 image with a fully-sampled T1 image by traditional method, illustrating the limitations of traditional method in this scenario; utilizing traditional methods to address the spatial alignment task and integrating it with the reconstruction task for joint optimization can further improve the spitial alignment performance as depicted in (c). However, the outcome is still inferior to that achieved using DUN-SA, as displayed in (d). Specifically, although the traditional method can achieve general alignment, as indicated by the blue arrow and yellow rectangle in the checkerboard patterns, it results in some local deformations due to the artifacts caused by undersampling of the target modality image and the inherent difficulties of the cross-modal spatial alignment task. These deformations are clearly displayed in the first row, highlighted by the yellow rectangle and green circle. In contrast, our method not only achieves general alignment but also preserves the local structures.
@@ 6.3 Comparison of spatial alignment performance
Additionally, we evaluate the spatial alignment performance of the proposed DUN-SA under different acceleration factors and compare it with SAN. In Fig. 11, (a) and (b) represent fully-sampled T2-weighted image and fully-sampled T1-weighted image, respectively. (c), (e), and (g) depict T1-weighted images aligned using SAN under different acceleration factors, while (d), (f), and (h) display T1-weighted images aligned using DUN-SA under different acceleration factors. We use a grid to facilitate the observation of spatial positions and employ a checkerboard to visualize misalignment. We find that both DUN-SA and SAN exhibit good spatial alignment performance at 4$\times$ acceleration. However, as the acceleration factor increases, SAN gradually fails to align T1 and T2 well, whereas DUN-SA is minimally affected. For instance, we can observe in the zoomed-in views of the third and fifth rows, at the locations indicated by the red, blue, and green arrows, misalignment still exists in (e) and (g), but this misalignment has been alleviated in (f) and (h). This demonstrates that DUN-SA, by iteratively solving the proposed model, appropriately integrates the spatial alignment task into the reconstruction process and achieves more precise spatial alignment results under high acceleration factors.
6.4 Adaption to scenarios with imperfect reference data
Table 5: Evaluation of adaptation to scenarios with imperfect reference data. US-Half/Rec-Half (US-All/Rec-all) indicates that half (all) reference modality data to be under-sampled/reconstructed.
Methods Equispaced 4$\times$ Equispaced 8$\times$
US-Half MC-CDic 37.28 $\pm$ 2.34 34.06 $\pm$ 3.09
0.9644 $\pm$ 0.0128 0.9398 $\pm$ 0.0293
MM-E2E-Varnet 38.73 $\pm$ 1.93 36.95 $\pm$ 2.07
0.9758 $\pm$ 0.0077 0.9674 $\pm$ 0.0106
SAN 38.64 $\pm$ 2.02 36.77 $\pm$ 2.22
0.9755 $\pm$ 0.0079 0.9663 $\pm$ 0.0115
DUN-SA 38.78 $\pm$ 2.03 37.02 $\pm$ 2.19
0.9760 $\pm$ 0.0080 0.9675 $\pm$ 0.0112
US-All MC-CDic 38.23 $\pm$ 1.66 35.66 $\pm$ 1.70
0.9719 $\pm$ 0.0077 0.9573 $\pm$ 0.0118
MM-E2E-Varnet 38.99 $\pm$ 1.62 36.79 $\pm$ 1.68
0.9767 $\pm$ 0.0066 0.9667 $\pm$ 0.0087
SAN 39.00 $\pm$ 1.62 37.04 $\pm$ 1.67
0.9762 $\pm$ 0.0065 0.9679 $\pm$ 0.0085
DUN-SA 39.87 $\pm$ 1.60 38.23 $\pm$ 1.66
0.9793 $\pm$ 0.0063 0.9724 $\pm$ 0.0082
Rec-Half MC-CDic 39.40 $\pm$ 1.79 36.88 $\pm$ 1.94
0.9772 $\pm$ 0.0073 0.9660 $\pm$ 0.0112
MM-E2E-Varnet 39.71 $\pm$ 1.80 37.79 $\pm$ 1.90
0.9791 $\pm$ 0.0068 0.9714 $\pm$ 0.0090
SAN 39.88 $\pm$ 1.79 38.21 $\pm$ 1.92
0.9796 $\pm$ 0.0066 0.9729 $\pm$ 0.0086
DUN-SA 40.74 $\pm$ 1.85 39.39 $\pm$ 2.02
0.9818 $\pm$ 0.0062 0.9771 $\pm$ 0.0079
Rec-All MC-CDic 38.88 $\pm$ 1.70 36.39 $\pm$ 1.80
0.9740 $\pm$ 0.0071 0.9630 $\pm$ 0.0104
MM-E2E-Varnet 39.22 $\pm$ 1.70 37.36 $\pm$ 1.77
0.9774 $\pm$ 0.0067 0.9693 $\pm$ 0.0087
SAN 39.32 $\pm$ 1.70 37.51 $\pm$ 1.81
0.9777 $\pm$ 0.0065 0.9699 $\pm$ 0.0085
DUN-SA 40.19 $\pm$ 1.73 38.64 $\pm$ 1.81
0.9803 $\pm$ 0.0061 0.9744 $\pm$ 0.0077
Fig. 12: The PSNR and SSIM curves on the fastMRI, IXI, In-house and BraTs datasets with different numbers of stages k.
Fig. 13: Visual comparison with effect of each component for 8$\times$ acceleration under 1D equispaced subsampling mask on the fastMRI dataset. First row: Reconstructed images by different methods; second row: Zoomed-in region of interest; third row: Equispaced mask of 8$\times$ acceleration and error maps of different methods.
In this section, we assess the adaptability of the reconstruction performance to scenarios with imperfect reference data. Specifically, during training, we set the reference modality (either all or half) to be under-sampled for both training and testing. From Table 5, we find that within the paired data, if the reference modality images are under-sampled, the reconstruction quality of the target modality images drops significantly. To mitigate this issue, we explore a method to enhance the flexibility and utility of our model. We train a single-modality reconstruction network (refer to ”w/o MM” in Section 6.6) for the reference modality, first reconstructing the under-sampled reference modality images and then using the reconstructed reference modality to assist in the reconstruction of the target modality. From the experimental results, all multi-modal methods have shown improvement. It is noteworthy that DUN-SA outperforms other multi-modal reconstruction methods across a range of reference data quality, from low to high. This demonstrates that DUN-SA also exhibits superior performance in scenarios with imperfect reference data.
@@ 6.5 Ablation Study
Table 6: Effect of each component on the performance of DUN-SA on the fastMRI, IXI, In-house and BraTs datasets for 4$\times$ and 8$\times$ acceleration under equispaced and random subsampling masks, measured in PSNR and SSIM.
Methods Equispaced 4$\times$ Equispaced 8$\times$ Random 4$\times$ Random 8$\times$
fastMRI w/o MM 39.36 $\pm$ 1.67 37.32 $\pm$ 1.76 43.35 $\pm$ 1.97 36.37 $\pm$ 1.77
0.9779 $\pm$ 0.0064 0.9694 $\pm$ 0.0082 0.9875 $\pm$ 0.0039 0.9645 $\pm$ 0.0096
w/o SA 41.07 $\pm$ 1.80 39.75 $\pm$ 1.95 44.68 $\pm$ 2.01 38.65 $\pm$ 1.98
0.9829 $\pm$ 0.0063 0.9787 $\pm$ 0.0078 0.9899 $\pm$ 0.0039 0.9755 $\pm$ 0.0090
w/o DB 41.16 $\pm$ 1.86 39.86 $\pm$ 1.88 44.70 $\pm$ 1.95 38.75 $\pm$ 2.02
0.9831 $\pm$ 0.0062 0.9791 $\pm$ 0.0075 0.9900 $\pm$ 0.0041 0.9758 $\pm$ 0.0092
DUN-SA 41.48 $\pm$ 1.84 40.23 $\pm$ 1.88 45.06 $\pm$ 1.85 39.23 $\pm$ 1.84
0.9838 $\pm$ 0.0060 0.9802 $\pm$ 0.0072 0.9907 $\pm$ 0.0029 0.9770 $\pm$ 0.0075
IXI w/o MM 42.54 $\pm$ 2.32 35.23 $\pm$ 2.09 40.36 $\pm$ 2.17 34.35 $\pm$ 2.05
0.9871 $\pm$ 0.0039 0.9571 $\pm$ 0.0099 0.9821 $\pm$ 0.0052 0.9501 $\pm$ 0.0107
w/o SA 47.12 $\pm$ 2.54 41.72 $\pm$ 2.33 45.27 $\pm$ 2.62 41.47 $\pm$ 2.32
0.9935 $\pm$ 0.0026 0.9848 $\pm$ 0.0053 0.9914 $\pm$ 0.0033 0.9843 $\pm$ 0.0055
w/o DB 46.94 $\pm$ 2.51 41.57 $\pm$ 2.31 44.98 $\pm$ 2.59 41.33 $\pm$ 2.30
0.9933 $\pm$ 0.0026 0.9846 $\pm$ 0.0053 0.9912 $\pm$ 0.0034 0.9839 $\pm$ 0.0056
DUN-SA 47.25 $\pm$ 2.53 41.84 $\pm$ 2.33 45.38 $\pm$ 2.51 41.62 $\pm$ 2.29
0.9936 $\pm$ 0.0026 0.9850 $\pm$ 0.0053 0.9915 $\pm$ 0.0032 0.9846 $\pm$ 0.0054
In-house w/o MM 35.74 $\pm$ 0.76 34.02 $\pm$ 0.77 37.78 $\pm$ 0.79 32.62 $\pm$ 0.74
0.9658 $\pm$ 0.0029 0.9531 $\pm$ 0.0039 0.9753 $\pm$ 0.0017 0.9458 $\pm$ 0.0042
w/o SA 37.06 $\pm$ 1.22 35.84 $\pm$ 1.31 40.02 $\pm$ 1.11 34.40 $\pm$ 1.34
0.9727 $\pm$ 0.0052 0.9663 $\pm$ 0.0069 0.9834 $\pm$ 0.0025 0.9563 $\pm$ 0.0082
w/o DB 37.20 $\pm$ 1.17 35.93 $\pm$ 1.37 39.94 $\pm$ 1.09 34.53 $\pm$ 1.16
0.9739 $\pm$ 0.0044 0.9671 $\pm$ 0.0067 0.9832 $\pm$ 0.0024 0.9584 $\pm$ 0.0061
DUN-SA 37.41 $\pm$ 1.08 36.31 $\pm$ 1.41 40.47 $\pm$ 1.09 34.84 $\pm$ 1.38
0.9747 $\pm$ 0.0039 0.9687 $\pm$ 0.0070 0.9845 $\pm$ 0.0024 0.9596 $\pm$ 0.0081
BraTs 2018 w/o MM 47.19 $\pm$ 2.77 37.55 $\pm$ 2.23 45.32 $\pm$ 2.38 37.16 $\pm$ 2.26
0.9934 $\pm$ 0.0026 0.9605 $\pm$ 0.0112 0.9921 $\pm$ 0.0030 0.9569 $\pm$ 0.0146
w/o SA 47.92 $\pm$ 2.83 39.16 $\pm$ 2.51 46.55 $\pm$ 2.51 38.91 $\pm$ 2.42
0.9947 $\pm$ 0.0024 0.9682 $\pm$ 0.0114 0.9930 $\pm$ 0.0028 0.9667 $\pm$ 0.0118
w/o DB 47.80 $\pm$ 2.78 38.98 $\pm$ 2.35 46.53 $\pm$ 2.53 38.74 $\pm$ 2.36
0.9945 $\pm$ 0.0022 0.9676 $\pm$ 0.0115 0.9929 $\pm$ 0.0028 0.9655 $\pm$ 0.0110
DUN-SA 48.14 $\pm$ 2.83 39.43 $\pm$ 2.40 46.93 $\pm$ 2.66 39.12 $\pm$ 2.42
0.9950 $\pm$ 0.0023 0.9694 $\pm$ 0.0111 0.9932 $\pm$ 0.0027 0.9682 $\pm$ 0.0114
To determine the optimal network architecture, we conduct two ablation studies. The first study focus on the number of stages, and the second emphasize each component within the network. The reference and target modalities are configured as follows: T1/T2 for fastMRI dataset; PD/T2 for IXI dataset; T1/T2 for In-house dataset; and T2/FLAIR for BraTs 2018 dataset.
@@ 6.5.1 Effect of number of stages
To demonstrate how the number of stages $k$ affects the reconstruction performance, we carry out a quantitative comparison of DUN-SA under different numbers of stages on four datasets mentioned for 4×4$\times$4 × and 8×8$\times$8 × acceleration under equispaced and random subsampling masks. Fig. 12 shows the mean SSIM and PSNR values for stages ranging from 1 to 16. We observe that from the first iteration to the twelfth iteration, the two metrics show a obviously progressive improvement. From the thirteenth iteration, the improvements in the two metrics became subtle or began to show a downward trend. Taking both performance and model complexity into consideration, we chose the model with $k=12$.
@@ 6.5.2 Effect of key components
Fig. 14: Visualization of Intermediate Results at Stage t: ground truth $x^{gt}$, reconstructed image $x^{t}$, warped reference image $\mathcal{T}(x_{\text{ref}},\phi^{t})$ denoted as $x_{ref}^{t}$, inter-modality prior $z^{t}$, and intra-modality prior $s^{t}$.
Based on the optimal number of iterations of 12 for the first ablation experiment, we conduct the second ablation experiment on our proposed DUN-SA on fastMRI, IXI, In-house, BraTs 2018 datasets as follows:
@@
Aligned Inter-modality Prior Learning Block (AIPLB): The ablation study for AIPLB means removing the AIPLB module from the network. It’s worth noting that, when the AIPLB module is removed, the assistance from the reference modality will also be omitted, and the spatial alignment operator will be simultaneously disregarded. The model degrades into a single-modal MRI reconstruction model. This configuration is termed the proposed DUN-SA without multi-modality (w/o MM).
Spatial Alignment Module (SAM): The ablation study for SAM involves removing the SAM module from the network. In this scenario, the spatial alignment operator is ignored within the network. As a result, the reference modality images are directly concated with intermediate reconstructed target modality images and feed in $\text{ProxNet}_{Z}$ to learn the inter-modality prior. This configuration is termed the proposed DUN-SA without spatial alignment (w/o SA).
@@
Denoising Block (DB): The ablation study for DB refers to the removal of the DB module from the network, which implies the elimination of the $\text{ProxNet}_{S}$’s function and the removal of intra-modality prior. This configuration is termed the proposed DUN-SA without denoising block (w/o DB).
Table 6 and Fig. 13 illustrate quantitative and qualitative comparisons of the reconstructed images from the four aforementioned methods in ablation study. From the results, we can find that (1) w/o MM shows the worst reconstruction results in terms of PSNR and SSIM. Additionally, from the error map, it is evident that this configuration results in the most severe loss of structural details and the most significant discrepancies compared to the ground-truth image. This clearly validates that the reference modality provides extra inter-modality prior to assist in the reconstruction of more structural details as presented in Section 3.1.1. (2) From the results, we can observe that w/o SA performs inferiorly compared to DUN-SA, as indicated by the brighter error map and lower metrics. This demonstrates that misalignment between modalities indeed affects the learning of inter-modality prior, as pointed out in Section 3.1.1, leading to $\text{ProxNet}_{Z}$’s insufficient exploration of the correlations between different modalities. Furthermore, the zoomed-in regions of interest shows that spatial alignment of corresponding points across modalities can further enable the network to more accurately capture structural information. For instance, as indicated by the arrows in the figure, DUN-SA can recover more clear details. (3) Compared to DUN-SA, w/o DB also exhibits inferior quantitative and qualitative reconstruction performance, indicating the effectiveness of intra-modality prior learning in DUN-SA.
Overall, each key component can help the model enhance reconstruction performance, and the optimal model can be obtained by combining all the modules.
6.6 Model verification
In this section, we utilize DUN-SA to conduct a model verification experiment to demonstrate the operational mechanism behind the network modules. Fig. 14 shows the reconstructed image $x^{t}$, warped reference image $\mathcal{T}(x_{\text{ref}},\phi^{t})$ denoted as $x_{ref}^{t}$, inter-modality prior $z^{t}$ and intra-modality prior $s^{t}$ at different stages. It can be easily observed that as the number of stages $t$ increases, the reference image gradually aligns with the fully sampled target image. The results verify the design of our optimization-inspired iterative learning framework—the mutual improvement of the spatial alignment task and the reconstruction task enables the proposed DUN-SA to achieve multi-modal MRI reconstruction as expected. Through these visualization results, the underlying rationale and insights of the proposed network can be intuitively understood, and our network demonstrates better transparency compared to other methods.
6.7 Computational Complexity Comparison
For the compared Multi-Modal MRI reconstruction methods (i.e., SAN, MC-CDic, and DUN-SA), Table 7 lists two important measures of the complexity, including the inference time and memory requirement for reconstruction of a slice with size 1 $\times$ 320 $\times$ 320 pixels and a 3D volume with size 16 $\times$ 320 $\times$ 320 pixels on an NVIDIA GeForce GTX 3090 GPU. As compared with other methods, DUN-SA exhibits mid-level complexity. However, considering the overall reconstruction performance, DUN-SA presents a favorable cost-performance trade-off.
Table 7: Complexity analysis of representative models.
Size Complexity SAN MC-CDic DUN-SA
1 Slice 320 $\times$ 320 Time (s) 0.50 2.31 1.09
Memory (MiB) 241.64 1089.90 398.66
16 Slices 320 $\times$ 320 Time (s) 2.51 33.02 7.02
Memory (MiB) 1338.79 16674.56 1554.41